OpenStudy (anonymous):

find dy/dx of y=(sqrt(11x) + (1/(sqrt(11))

OpenStudy (anonymous):

\[y=(\sqrt(11x) + (1/(\sqrt(11))\]

OpenStudy (turingtest):

\[y=(11x)^{1/2}+11^{-1/2}\] use the chain rule and realize the last term is a constant, and so will disappear upon taking the derivative: \[dy/dx=(1/2)(11x)^{-1/2}(11)=11/(2\sqrt{11x})\]